the lunch box is expanding and expects operating cash flows of 32500 a year for three years as a result. this expansion requires 28000 in new fixed assets. these assets will be worthless and fully depreciated at the end of the project. in addition, the project requires an initial investment of 2800 in net working capital at the beginning of the project, which will be recovered at the end of the project. what is the net present value of this expansion project at a required rate of return of 16 percent?
Answer #1

Initial investment = Cost + increase in NWC

Initial investment = 28,000 + 2,800 = 30,800

Year 3 total cash flow = 2,800 + 32500 = 35,300

Net present value = Present value of cash inflows - present value of cash outflows

Net present value = -30,800 + 32500 / (1 + 0.16)1 + 32500 / (1 + 0.16)2 + 35,300 / (1 + 0.16)3

Net present value = -30,800 + 28,017.2414 + 24,152.7943 + 22,615.2159

Net present value = $43,985.25